# Galaxy Morphologies at Cosmic Noon with JWST: A Foundation for Exploring Gas Transport with Bars and Spiral Arms

### v1

@@ 2.3 Kinematics sample
To trace the dynamical evolution of galaxies at the epochs probed in this study, we also examine the interplay between morphological features observed in the JWST NIRCam images and available kinematic measurements, specifically the $v/\sigma$ ratio, which provides crucial insights into each galaxy’s dynamical state. To compile the kinematic data, we use measurements from well-established surveys and studies.
@@
We found an overlap of 155 galaxies with existing measurements of velocity and velocity dispersion. For consistency, we adopt a standardized approach, using the rotation velocity at one effective radius, $v_{\mathrm{rot}}(r_{\mathrm{eff}})$, and assumed a radially constant intrinsic disk velocity dispersion $\sigma_{0}$, corrected for observational effects, such as beam smearing, as derived from modelling. The detailed measurement adjustments are outlined as follows:
@@
  • In Wisnioski et al. (2019), from a total number of 745, we identified 7 more overlapping galaxies.
  • Lastly, for Price et al. (2020) (MOSDEF), we identified 26 overlapping galaxies from their full sample of 1,717, of which we used 23. Since these measurements were based on slit spectroscopy with the MOSFIRE instrument, the slits could not be precisely aligned with the galaxies’ major axes. The sample of 23 used here is restricted to objects that are spatially resolved (under seeing-limited conditions), with reasonable alignment between slit and galaxy major axes ($|\Delta\mathrm{PA}|\leq 45^{\circ}$), and have detected rotation signatures (within the uncertainties). We note that this work used an arctan parametric fit to derive the rotation velocity profile, differing from the parametric fits used in other surveys. Moreover, as the exact kinematic major axes were unknown, the modelling assumed the morphological and kinematic axes were the same, introducing potential systematics. However, given the small sample size, any such effect is unlikely to have a significant impact on our overall analysis.
@@
Table 2: Classification results based on our 4/6 expert vote threshold, showing the main morphological categories and their distribution across redshift bins. The top rows list total fractions normalized by the full sample, while the middle section details subcategories among disk galaxies. The bottom rows present the distribution of spiral arm counts.
@@
Disk-like $N_{\rm disk}/N_{\rm total}$ 1,186 82 727 90 459 72
Compact/Sph $N_{\rm c,s}/N_{\rm total}$ 24 2 17 2 7 1
Int/Merger $N_{\rm i,m}/N_{\rm total}$ 139 10 48 6 91 14
Undefined $N_{\rm undefined}/N_{\rm total}$ 22 2 11 1 11 2
+ Lens systems $N_{\rm lens}/N_{\rm total}$ 3 $<1$ 1 1 2 $<1$
+ Companion $N_{\rm+\,companion}/N_{\rm total}$ 67 5 32 4 35 6
No 4/6 agreement $N_{\rm no-agree}/N_{\rm total}$ 80 6 8 1 72 11
Spirals $N_{\rm spiral}/N_{\rm total}$ 574 40 444 55 130 20
Bars $N_{\rm bar}/N_{\rm total}$ 43 3 42 5 1 $<1$
Bulges $N_{\rm bulge}/N_{\rm total}$ 703 48 495 61 208 33
Spirals $N_{\rm spiral}/N_{\rm disk}$ 574 48 444 61 130 28
Bars $N_{\rm bar}/N_{\rm disk}$ 43 4 42 6 1 $<1$
Bars (isophotal fitting) $N_{\rm bar}^{(a)}/N_{\rm disk}^{(b)}$ 84 11 74 16 10 3
Bulges $N_{\rm bulge}/N_{\rm disk}$ 703 59 495 68 208 45
Clumpy $N_{\rm clumpy}/N_{\rm disk}$ 140 12 81 11 59 13
Edge-on $N_{\rm edge-on}/N_{\rm disk}$ 167 14 113 16 54 12
Two-arm spirals $N_{\rm 2,arm}/N_{\rm spiral}^{(c)}$ 208 61 156 60 52 64
Three-arm spirals $N_{\rm 3,arm}/N_{\rm spiral}^{(c)}$ 104 30 79 30 25 31
Four-arm spirals $N_{\rm 4,arm}/N_{\rm spiral}^{(c)}$ 18 5 15 6 3 4
777 \$a\$\$a\$footnotetext: These values correspond to the classification using isophotal fitting on a larger candidate sample, selected with a 2/6 voting threshold and further assessed as detailed in §3.2.
@@
\$c\$\$c\$footnotetext: $N_{\mathrm{spiral}}$ in the bottom rows corresponds to the number of galaxies for which a 4/6 agreement was reached (343 galaxies).
Figure 6: Global classification results for the 1,451 galaxies in our sample. The blue and orange bars represent the low- and high-redshift bins, respectively, while the white bars correspond to the full sample. The percentages on top of the bars correspond to fractions normalized by the number of galaxies in the full sample, with black for the total, and blue and orange for the two redshift bins. The inset figure highlights the galaxies classified as disks in the sample, for which we also note the presence or not of five morphological features. The percentages displayed on top of those bars indicate the fractions of each subcategory normalized by the total number of disks. The $\dagger$ symbol in “+ Companion” denotes a non-mutually exclusive category, meaning it can be assigned alongside any of the main categories.
@@ 3.1 Classification scheme
To prepare the JWST NIRCam imaging for analysis, we first fit the images with a single Sérsic profile using galfit (Peng et al., 2002, 2010) for each stamp. Initial source detection relied on segmentation maps, which provided preliminary inputs for the galfit routine. Following this, six of the authors of this work (J.M.E.S, C.B, S.P, C.P, J.C, J-B.J) visually classified all the galaxies in the sample. For increased robustness, the morphological classification process incorporated available image stamps, colour composites, residual images, and the best-fit Sérsic indices. Additionally, unsharp-masked versions of the images were also shown to the classifiers to enhance fine details and highlight distinct features that might otherwise be overlooked. To facilitate the classification process, given the large sample size of approximately 10,000 images, we used a custom classification tool for galaxy morphologies (galclass  [^8] https://github.com/spastras/galclass).
@@
i) Disk: These galaxies are characterized by a distinct, organized structure with a Sérsic index close to unity, consistent with Wuyts et al. (2011b), who showed that $n\sim 1$ is typical for main sequence galaxies. They often show spiral arms and may have a central bulge. Depending on the viewing angle, they can appear circular, oval, or linear, corresponding to face-on, intermediate, or edge-on orientations.
ii) Compact-Spheroidal: These systems appear smooth and rounded, with no visible disk or spiral structure. They have a uniform brightness profile and are typically featureless. When classifying these galaxies, we also considered their Sérsic indices, selecting systems that were well described by $n\approx 4$.
@@
Within each category, we also included an option to identify the presence of companion galaxies as well as a strongly lensed system, identified by the presence of arcs, or rings of a background object gravitationally lensed by the foreground galaxy. For galaxies classified as disks, we also noted the presence or not of five non-exclusive morphological features labelled “Spiral”, “Bar”, “Bulge”, “Clumpy”, and “Edge-on”, several of which may be present in a given galaxy. Figure 4 provides an overview of the classification scheme, showcasing the main categories and representative examples of each category and subcategory with colour composite images. These images were generated using trilogy (Coe et al., 2012), with filter groupings carefully selected to enhance the visibility of key galaxy features. Specifically, shorter-wavelength filters were assigned to blue, longer-wavelength filters to red, and intermediate-wavelength filters to green, ensuring a balanced colour representation. An equal number of filters (where possible) were distributed across the RGB channels to maintain consistency in the visualization.  [^9] The individual images used to generate the colour composites were not PSF-matched. While variations in PSF size exist across filters, the impact on the colour images is generally minor, as the broad structural features of galaxies remain visually consistent. Moreover, these images are intended solely for visual representation of the sample, and no quantitative measurements are derived from them.
The final classification was determined based on a consensus threshold, requiring at least 4 out of 6 expert votes for a category assignment. We emphasize that galaxies failing to meet this threshold for any category remain unclassified, which is distinct from the “Undefined” category, as the latter represents galaxies with inherently disordered shapes or non-coherent structures, and is thus a category on its own.
@@ 3.2 Classification overview
Our visual classification results, based on the 4/6 voting threshold (Table 2 and Figure 6) reveal that the majority of galaxies in our sample have a disk-like morphology, comprising 82% of the total sample ($\approx 90\%$ and $\approx 72\%$ in the low- and high-redshift bins, respectively). Interacting/merging galaxies form the second-largest category, accounting for $\approx 10\%$ of the sample, with a notable increase in the higher redshift bin. Spheroidal galaxies make up only 1.7% of the sample, with slightly more found in the lower redshift bin. Lensed systems and undefined galaxies are rare, representing 0.1% and 1.5%, respectively. Notably, only 80 galaxies (5.5%) lacked a 4/6 voting agreement.
Figure 7: Disk fraction as a function of effective radius ($r_{\mathrm{eff}}$), measured in bins of $\Delta r=0.1$ arcsec. The blue points represent the disk fraction in each bin, with error bars indicating binomial uncertainties. The dashed line connects the points for visualization. The fraction steadily increases with effective radius, which may reflect an intrinsic trend or the greater ease of visually identifying disks in larger galaxies. The histogram in the top panel shows the number of galaxies per bin, with lower counts in the first and last bins (5 and 8 galaxies, respectively), contributing to the larger uncertainties in those bins.
An important potential bias to consider in the classification of galaxies (whether through visual or automatic methods) is related to their apparent size. Specifically, we observe increasing disk fractions $f_{\mathrm{disk}}$ as a function of $r_{\mathrm{eff}}$ as shown in Figure 7, where we show $f_{\mathrm{disk}}$ in bins of $\Delta r=0.1$ arcsec. This is expected since it is easier to identify a disk-like structure in extended sources. However, this trend may also have a physical origin, as disks typically exhibit extended light distributions, which naturally results in larger effective radii and contributes to the observed correlation.
A more extended discussion of observational limitations, such as resolution differences and surface brightness dimming effects, is provided in Appendix A.
@@ 3.3.1 Spiral structure
Beyond their presence alone, the number of spiral arms also encodes information about the underlying dynamics (e.g., Elmegreen & Elmegreen, 2014; Dobbs & Baba, 2014). While two-armed (grand-design) spirals are likely driven by global disk modes or external perturbations, such as bars or tidal interactions (e.g., Toomre & Toomre, 1972; Athanassoula, 1984; Dobbs et al., 2010; D’Onghia et al., 2013; Kendall et al., 2015), multi-armed (flocculent) spirals typically emerge in marginally stable disks through mechanisms like swing amplification (Toomre, 1981). As such, both the detection and characterization of spiral arms (particularly arm multiplicity) offer key diagnostics of disk structure, stability, and the mechanisms shaping galaxy evolution (e.g., Hart et al., 2017).
To investigate potential trends in spiral structure across our two redshift bins, we visually counted the number of arms within the spiral galaxies identified in our classification (574 in total). Following a similar approach used for morphological classification, six authors (J.M.E.S., C.B., C.P., J.C., J-B.J., G.T.) independently performed the arm counts. To aid in the identification of spiral arms, we utilized a custom interactive tool for visualizing FITS files with customizable scales using all the stamps for each galaxy. Each team member independently determined the number of visible arms for each galaxy. Figure 5 (bottom panels) shows examples of four galaxies with a consensus on the number of counted spiral arms. Figure 5 (bottom panels) provides examples of galaxies with consensus arm counts, while full-colour composite images of all classified spiral galaxies are available at this link. The results of the spiral arm counting are discussed in §4.
@@ 3.3.2 Bar structure
Figure 8: Bar identification analysis for GN4_10280: Left: F [^4] W image with the best-fit major axis of the bar indicated by the dashed line. Middle: Radial profiles of ellipticity ($e$, top) and position angle (PA, bottom) as a function of the major axis of the isophotes, extracted from the F [^4] W image. The vertical dashed line marks the maximum in the ellipticity. Right: RGB colour composite using F182M (blue), F210M (green), and F [^4] W (red), highlighting the galaxy’s structure in different wavelengths.
We identified barred galaxies in our sample by refining the initial visual classification through the fitting of elliptical isophotes. Since visually identifying bars can be challenging as they can be obscured by bright central bulges or may blend with spiral arms, we applied a more inclusive threshold of 2/6 classifier votes to ensure potential barred galaxies were not overlooked. The isophotal fitting method provides a quantitative test for confirming bar structures.
Starting from a parent sample of 123 galaxies that received at least 2/6 classifier votes for a bar, we applied an ellipse-fitting technique using photutils (Bradley et al., 2016) on images taken with the longest available wavelength filter (F [^4] W, except for one case where F356W was used). We then analyzed the resulting ellipticity ($e=1-q$), where $q$ is the axis ratio, as well as the position angle profiles as functions of radius to confirm the presence of bars. To classify a galaxy as barred, we imposed the following criteria: i) the galaxy must not be edge-on, defined as having an ellipticity below 0.5 at large radii, outside the regions influenced by the bar or spiral arms; ii) The ellipticity must increase smoothly with radius, reaching a well-defined peak above 0.25 before decreasing again, as expected for a bar. This ensures that the feature is significantly elongated and not an artifact of noise or a mild bulge elongation; and iii) the position angle must remain relatively constant (within 20°). We show an example of this analysis in Figure 8, with the identification of a bar for galaxy GN4_10280.
With these additional criteria, the number of identified barred systems increased to 84, with 74 in the low-redshift bin and 10 in the high-redshift bin. Notably, 81 out of the 84 barred galaxies were also classified as spiral galaxies. Figure 5 (top panels) presents four examples of barred galaxies in our sample. The full mosaic of the colour images of the barred systems is available at this link [^10] https://www.mpe.mpg.de/resources/IR/JWSTColorImages/barred_grid_A4.pdf.
[…]

### v2

@@ 2.3 Kinematics sample
To trace the dynamical evolution of galaxies at the epochs probed in this study, we also examined the interplay between morphological features observed in the JWST NIRCam images and available kinematic measurements, specifically the $v/\sigma$ ratio, which provides crucial insights into each galaxy’s dynamical state. To compile the kinematic data, we used measurements from well-established surveys and studies.
@@
We found an overlap of 155 galaxies that have velocity and velocity dispersion measurements. For consistency, we adopted a standardized approach, using the rotation velocity at one effective radius, $v_{\mathrm{rot}}(r_{\mathrm{eff}})$, and assume a radially constant intrinsic disk velocity dispersion $\sigma_{0}$, corrected for observational effects, such as beam smearing, as derived from modeling. The detailed measurement adjustments are outlined as follows:
@@
  • In Wisnioski et al. (2019), from a total number of 745, we identified seven more overlapping galaxies.
  • Lastly, for Price et al. (2020) (MOSDEF), we identified 26 overlapping galaxies from their full sample of 1,717, of which we used 23. Since these measurements were based on slit spectroscopy from the MOSFIRE instrument, the slits could not be precisely aligned with the galaxies’ major axes. The sample of 23 used here is restricted to objects that are spatially resolved (under seeing-limited conditions), have reasonable alignment between slit and galaxy major axes ($|\Delta\mathrm{PA}|\leq 45^{\circ}$), and have detected rotation signatures (within the uncertainties). We note that this work used an arctan parametric fit to derive the rotation velocity profile, differing from the parametric fits used in other surveys. Moreover, as the exact kinematic major axes were unknown, the modeling assumed the morphological and kinematic axes were the same, introducing potential systematics. However, given the small sample size, any such effect is unlikely to have a significant impact on our overall analysis.
@@
Figure 7: Global classification results for the 1,451 galaxies in our sample. The blue and orange bars represent the low- and high-redshift bins, respectively, while the white bars correspond to the full sample. The percentages on top of the bars correspond to fractions normalized by the number of galaxies in the full sample, where black indicates the total, and blue and orange the two redshift bins. While the main text reports fractions with associated uncertainties, we show percentages in this figure to facilitate readability and visual comparison across bins. The inset figure highlights the galaxies classified as disks in the sample, for which we also note the presence of five morphological features. The percentages displayed on top of those bars indicate the fractions of each subcategory normalized by the total number of disks. The $\dagger$ symbol in + Companion denotes a nonmutually exclusive category, meaning it can be assigned alongside any of the main categories.
Table 2: Classification results overview
@@
Disk-like $N_{\rm disk}/N_{\rm total}$ 1,186 $82\pm 3$ 727 $90\pm 3$ 459 $72\pm 5$
Compact/Sph $N_{\rm c,s}/N_{\rm total}$ 24 $2\pm 1$ 17 $2\pm 2$ 7 $1\pm 1$
Int/Merger $N_{\rm i,m}/N_{\rm total}$ 139 $10\pm 2$ 48 $6\pm 2$ 91 $14\pm 4$
Undefined $N_{\rm undefined}/N_{\rm total}$ 22 $2\pm 1$ 11 $1\pm 1$ 11 $2\pm 2$
+ Lens systems $N_{\rm lens}/N_{\rm total}$ 3 $<1\pm 0$ 1 $1\pm 0$ 2 $<1\pm 1$
+ Companion $N_{\rm+\,companion}/N_{\rm total}$ 67 $5\pm 2$ 32 $4\pm 2$ 35 $6\pm 3$
No 4/6 agreement $N_{\rm no-agree}/N_{\rm total}$ 80 $6\pm 2$ 8 $1\pm 1$ 72 $11\pm 4$
Spirals $N_{\rm spiral}/N_{\rm total}$ 574 $40\pm 4$ 444 $55\pm 5$ 130 $20\pm 5$
Bars $N_{\rm bar}/N_{\rm total}$ 43 $3\pm 1$ 42 $5\pm 2$ 1 $<1\pm 0$
Bulges $N_{\rm bulge}/N_{\rm total}$ 703 $48\pm 4$ 495 $61\pm 5$ 208 $33\pm 6$
Spirals $N_{\rm spiral}/N_{\rm disk}$ 574 $48\pm 4$ 444 $61\pm 5$ 130 $28\pm 6$
Bars $N_{\rm bar}/N_{\rm disk}$ 43 $4\pm 2$ 42 $6\pm 3$ 1 $<1\pm 1$
Bars (isophotal fitting) $N_{\rm bar}^{(a)}/N_{\rm disk}^{(b)}$ 84 $11\pm 3$ 74 $16\pm 5$ 10 $3\pm 3$
Bulges $N_{\rm bulge}/N_{\rm disk}$ 703 $59\pm 4$ 495 $68\pm 5$ 208 $45\pm 7$
Clumpy $N_{\rm clumpy}/N_{\rm disk}$ 140 $12\pm 3$ 81 $11\pm 4$ 59 $13\pm 5$
Edge-on $N_{\rm edge-on}/N_{\rm disk}$ 167 $14\pm 3$ 113 $16\pm 4$ 54 $12\pm 5$
Two-arm spirals $N_{\rm 2,arm}/N_{\rm spiral}^{(c)}$ 208 $61\pm 8$ 156 $60\pm 9$ 52 $64\pm 16$
Three-arm spirals $N_{\rm 3,arm}/N_{\rm spiral}^{(c)}$ 104 $30\pm 7$ 79 $30\pm 9$ 25 $31\pm 15$
Four-arm spirals $N_{\rm 4,arm}/N_{\rm spiral}^{(c)}$ 18 $5\pm 4$ 15 $6\pm 4$ 8 $4\pm 6$
777 All the fractions above are based on our 4/6 expert vote threshold. We show the main morphological categories with their distribution (in percentage) across redshift bins and corresponding $3\sigma$ standard errors. The top rows list total fractions normalized by the full sample, while the middle section details subcategories among disk galaxies. The bottom rows present the distribution of spiral arm counts.
\$a\$\$a\$footnotetext: These values correspond to the classification using isophotal fitting on a larger candidate sample, selected with a 2/6 voting threshold and further assessed as detailed in Sect. 3.2.
@@
\$c\$\$c\$footnotetext: $N_{\mathrm{spiral}}$ in the bottom rows corresponds to the number of galaxies for which a 4/6 agreement was reached (341 galaxies).
@@ 3.1 Classification scheme
To prepare the JWST NIRCam imaging for analysis, we first fitted the images using a single Sérsic profile with galfit (Peng et al., 2002, 2010) for each stamp. Initial source detection relied on segmentation maps, which provided preliminary inputs for the galfit routine. Following this, six of the authors of this work (J.M.E.S, C.B, S.P, C.P, J.C, J-B.J) visually classified all the galaxies in the sample. For increased robustness, the morphological classification process incorporated available image stamps, color composites, residual images, and the best-fit Sérsic indices. Additionally, unsharp-masked versions of the images were also shown to the classifiers to enhance fine details and highlight distinct features that might otherwise be overlooked. To facilitate the classification process, given the large sample size of approximately 10,000 images, we used a custom classification tool for galaxy morphologies (GalClass; Pastras, 2025) [^8] Available at https://github.com/spastras/galclass.
@@
i) Disk: These galaxies are characterized by a distinct, organized structure with a Sérsic index close to unity, consistent with Wuyts et al. (2011b), who showed that $n\sim 1$ is typical for main-sequence galaxies. They often show spiral arms and may have a central bulge. Depending on the viewing angle, they can appear circular, oval, or linear, corresponding to face-on, intermediate, or edge-on orientations.
ii) Compact-Spheroidal: These systems appear smooth and rounded, and show no visible disk or spiral structure. They have a uniform brightness profile and are typically featureless. When classifying these galaxies, we also considered their Sérsic indices, selecting systems that were well described by $n\approx 4$.
@@
Within each category, we also included an option to identify the presence of companion galaxies as well as a strongly lensed system, identified by the presence of arcs, or rings of a background object gravitationally lensed by the foreground galaxy. For galaxies classified as disks, we also noted the presence or not of five non-exclusive morphological features labeled Spiral, Bar, Bulge, Clumpy, and Edge-on, several of which may be present in a given galaxy. Figure 4 shows an example of GalClass for a spiral galaxy, and Figure 5 provides an overview of the classification scheme, showcasing the main categories and representative examples of each category and subcategory with color composite images. These images were generated using trilogy (Coe et al., 2012), using filter groupings carefully selected to enhance the visibility of key galaxy features. Specifically, shorter-wavelength filters were assigned to blue, longer-wavelength filters to red, and intermediate-wavelength filters to green, ensuring a balanced color representation. An equal number of filters (where possible) were distributed across the RGB channels to maintain consistency in the visualization.  [^9] The individual images used to generate the color composites were not PSF-matched. While variations in PSF size exist across filters, the impact on the color images is generally minor, as the broad structural features of galaxies remain visually consistent. Moreover, these images are intended solely for visual representation of the sample, and no quantitative measurements are derived from them.
The final classification was determined based on a consensus threshold, requiring at least four out of six expert votes for a category assignment. We emphasize that galaxies failing to meet this threshold for any category remain unclassified, which is distinct from the “Undefined” category, as the latter represents galaxies that have inherently disordered shapes or noncoherent structures, and is thus a category on its own. Figure 6 illustrates representative cases of galaxies identified as bars and spirals, each meeting the consensus threshold of four out of six expert votes.
[…]
@@ 3.2 Classification overview
Our visual classification results, based on the 4/6 voting threshold (Table 2 and Figure 7), reveal that the majority of galaxies in our sample have a disk-like morphology, comprising a fraction of $0.82\pm 0.03$ of the total sample ($0.90\pm 0.03$ and $0.72\pm 0.05$ in the low- and high-redshift bins, respectively). Interacting or merging galaxies form the second-largest category, accounting for approximately $0.10\pm 0.02$ of the sample, and showing a notable increase in the higher-redshift bin. Spheroidal galaxies make up a small fraction ($0.02\pm 0.01$), with slightly more found in the lower-redshift bin. Lensed systems and undefined galaxies are rare, each contributing less than 0.01. Notably, only 80 galaxies (fraction $\approx 0.06\pm 0.02$) lacked a 4/6 voting agreement.
Figure 8: Disk fraction as a function of effective radius $r_{\mathrm{eff}}$, measured in angular units with bin size $\Delta r=0.1$ arcsec (top) and in physical units of $\Delta r=1$ kpc (bottom). Blue points represent the disk fraction in each bin, and error bars indicate binomial uncertainties. Dashed lines are shown to guide the eye. In both panels, the disk fraction increases with effective radius, highlighting that disks are more easily identified in galaxies with larger apparent and physical sizes.
Figure 9: Bar identification analysis for two example galaxies. Top: GN4_10280, identified as barred based on both visual morphology (N=6 votes) and quantitative isophotal fitting. Left panel: F [^4] W image with best-fit bar major axis (dashed). Middle: radial profiles of ellipticity ($e$) and position angle (PA) with the ellipticity peak marked. Right: RGB composite showing structural features. Bottom: COS4_01493, initially not classified as barred via visual inspection (N=3 votes), but later confirmed as barred through isophotal fitting.
An important potential bias to consider in the classification of galaxies (whether through visual or automatic methods) is related to their apparent size. Specifically, we observe increasing disk fractions $f_{\mathrm{disk}}$ as a function of size as shown in Figure 8, where we show $f_{\mathrm{disk}}$ as a function of effective radius in apparent size (arcsec) and physical size (kpc). This is expected since it is easier to identify a disk-like structure in extended sources. However, this trend may also have a physical origin, as disks typically exhibit extended light distributions, which naturally result in larger effective radii and contribute to the observed correlation. An additional discussion on observational limitations, such as spatial resolution, surface brightness dimming, and wavelength coverage, is provided in Appendix A.
@@ 3.3.1 Spiral structure
Beyond their presence alone, the number of spiral arms also encodes information about the underlying dynamics (e.g., Elmegreen & Elmegreen, 2014; Dobbs & Baba, 2014). While two-armed (grand-design) spirals are likely driven by global disk modes or external perturbations, such as bars or tidal interactions (e.g., Toomre & Toomre, 1972; Athanassoula, 1984; Dobbs et al., 2010; D’Onghia et al., 2013; Kendall et al., 2015), multi-armed (flocculent) spirals typically emerge in marginally stable disks through mechanisms such as swing amplification (Toomre, 1981). As such, both the detection and characterization of spiral arms (particularly arm multiplicity) offer key diagnostics of disk structure, stability, and the mechanisms shaping galaxy evolution (e.g., Hart et al., 2017).
To investigate potential trends in spiral structure across our two redshift bins, we visually counted the number of arms within the spiral galaxies identified in our classification (574 in total). Following a similar approach used for morphological classification, six authors (J.M.E.S., C.B., C.P., J.C., J-B.J., G.T.) independently performed the arm counts. To aid in the identification of spiral arms, we utilized a custom interactive tool for visualizing FITS files with customizable scales using all the stamps for each galaxy. Each team member independently determined the number of visible arms for each galaxy. Figure 6 (bottom panels) provides examples of galaxies with consensus arm counts, while full-color composite images of all classified spiral galaxies are available at this link. The results of the spiral arm counting are discussed in Sect. 4.
@@ 3.3.2 Bar structure
We identified barred galaxies in our sample by refining the initial visual classification through the fitting of elliptical isophotes. Visually identifying bars can be challenging as they can be obscured by bright central bulges or may blend with spiral arms. Therefore, we applied a more inclusive threshold of 2/6 classifier votes to ensure potential barred galaxies were not overlooked. The isophotal fitting method provides a quantitative test for confirming bar structures.
Starting from a parent sample of 123 galaxies that received at least 2/6 classifier votes for a bar, we applied an ellipse-fitting technique using photutils (Bradley et al., 2016) on images taken with the longest available wavelength filter (F [^4] W, except for one case where F356W was used). We then analyzed the resulting ellipticity ($e=1-q$), where $q$ is the axis ratio, as well as the position angle profiles as functions of radius to confirm the presence of bars. To classify a galaxy as barred, we imposed the following criteria: i) the galaxy must not be edge-on, defined as having an ellipticity below 0.5 at large radii, outside the regions influenced by the bar or spiral arms; ii) The ellipticity must increase smoothly with radius, reaching a well-defined peak above 0.25 before decreasing again, as expected for a bar. This ensures that the feature is significantly elongated and not an artifact of noise or a mild bulge elongation; and iii) the position angle must remain relatively constant (within 20°). We show two examples of this analysis in Figure 9: galaxy GN4_10280, which received N=6 votes for barred system, and galaxy COS4_01930, which only received N=3 votes but was later confirmed barred following this quantitative analysis.
Using these additional criteria, the number of identified barred systems increased to 84, where 74 are in the low-redshift bin and ten are in the high-redshift bin. Notably, 81 out of the 84 barred galaxies were also classified as spiral galaxies. Figure 6 (top panels) presents four examples of barred galaxies in our sample. The full mosaic of the color images of the barred systems is available at this link [^10] https://www.mpe.mpg.de/resources/IR/JWSTColorImages/barred_grid_A4.pdf.